FUTURE VALUE PROBLEMS he Be year of nevolent Company has agreed to lend you funds to complete the last your degree. The Company will lend you $2,400 today, if you agree to y a lump sum of $4,000 4 years from now. What is the approximate annual rate of interest that Benevolent is charging you? You discover $40,000 under your pillow, which can be invested at a rate of 8% 2. per year, compounded semi-annually. If you spend $11,435 per year, how long will the money last? What amount would you have in a retirement account if you made monthly deposits of $1000 for 20 years earning 9%, compounded quarterly? 3. PRESENT VALUE PROBLEMS 1. Your parents have promised to give you a graduation present of $5,000 when you graduate in four years time. If interest rates stay at 6% compounded annually for the next four years, how much is this money worth in todays dollars? 4 6 2. To have $6,000 for a childs education in 10 years what amount should a parent deposit in a savings account now that earns 12%, compounded quarterly? CIN

Please full calculation and explanation.
Answer #1

1. Principle=2400

Time = 4 years

sum= 4000

Total interest= sum - principal= 4000-2400= 1600

simple interest=( principle× rate×time)/100

1600= (2400×r×4)/100= 16.67%(rate of interest)

2.40,000 = $11,435 (PVAF 18%, n)

= PVAF 18%, n)

Solution n= 6 years.

Pesent value problems

1.Amount invested today = FV = $5000

Return expected from investment = i= 6

Duration of investment = n= 4 years

Frequency of compounding = m= 1

FV= pv( 1+r/100)n = 5000= PV( 1+ o.o6)4

PV=3968

2. Future value= 6000

time(t)= 10 years

rate of interest(r)=12

Future value = present value(1+r)t

6000=pv(1+0.12)10

solving we get =1935

Image of page 6